OpenStudy (anonymous):

Will Medal And Fan Best Answer!! I need explanation AND answer! Write 43/25 as a percentage.

OpenStudy (anonymous):

1. Divide 43/25. 2. Multiply the quotient by 100.

OpenStudy (anonymous):

Use your calculator. \[\frac{ 43 }{ 25 }\times100\]

OpenStudy (anonymous):

1.72 doesn't help

OpenStudy (anonymous):

whats with the decimal, its obiously not what im looking for

OpenStudy (anonymous):

That's why you multiply it by 100.

OpenStudy (anonymous):

Read and understand.

OpenStudy (anonymous):

Expressed as a decimal, 43/25 is equal to 1.72.

OpenStudy (anonymous):

take away the decimal

OpenStudy (anonymous):

you end up with?

OpenStudy (anonymous):

172

OpenStudy (anonymous):

172%

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

yes good job

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

can you answer one more?lol

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Write 48/50 as a percentage.

OpenStudy (anonymous):

= 48/50 * 100% = 0.96 * 100% = 96%

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Np